# Neural Approximate Mirror Maps for Constrained Diffusion Models

### v1

@@ 3.2 Learning the mirror diffusion model
Similarly to Liu et al. [47], we train an MDM on the mirror distribution ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$ and map its samples to the constrained space through $\mathbf{f}_{\psi}$. In particular, we train a score model $\mathbf{s}_{\theta}$ with the following denoising score matching objective in the learned mirror space (defined as the range of $\mathbf{g}_{\phi}$):
@@ 3.3 Finetuning the inverse mirror map
The inverse map $\mathbf{f}_{\psi}$ is trained with samples from the noisy mirror distributions in Eq. 2, but we ultimately wish to evaluate $\mathbf{f}_{\psi}$ with samples from the MDM. To reduce the distribution shift, it may be helpful to finetune $\mathbf{f}_{\psi}$ with MDM samples. In the following finetuning objective, we replace $\tilde{\mathbf{x}}\sim{(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$ with $\tilde{\mathbf{x}}\sim p_{\theta}$, where $p_{\theta}$ is the distribution of MDM samples in the mirror space:
@@
Finetuning essentially tailors $\mathbf{f}_{\psi}$ to the MDM. The original objective assumes that the MDM will sample Gaussian-perturbed images from the mirror distribution, but in reality it samples from a slightly different distribution. We generate a training dataset of samples $\tilde{\mathbf{x}}$ from the MDM and then finetune the inverse map to deal with such samples specifically.
@@ 4 Results
We present experiments with constraints ranging from physics-based to semantic, finding improved constraint satisfaction with greater training efficiency. Appendices A.1 and B provide implementation and constraint details, respectively. The following paragraphs introduce the demonstrated constraints $\ell$. For each constraint, we consider an image dataset for which the constraint is physically meaningful.
@@
In astronomical imaging, even if the structure of a source is unknown a priori, often the total brightness, or total flux, is well constrained [24, 25]. We define $\ell_{\text{flux}}(\mathbf{x})$ as the absolute difference between the sum of the pixel values in $\mathbf{x}$ and the true total brightness. For demonstration, we use a dataset of $64\times 64$ images of black-hole simulations [68] whose pixel values sum to $120$.
@@
We consider Burgers’ equation [7, 13] for a 1D viscous fluid, representing the discretized solution as an $n_{x}\times n_{t}$ image $\mathbf{x}$, where $n_{x}$ and $n_{t}$ are the numbers of grid points in space and time, respectively. The distance $\ell_{\text{burgers}}(\mathbf{x})$ compares each 1D state in the image to the PDE solver’s output given the previous state (based on Crank-Nicolson time-discretization [19, 41]). The dataset consists of $64\times 64$ images with Gaussian random fields as initial conditions.
@@
A time-dependent 2D velocity field $\mathbf{u}=\mathbf{u}(x,y,t)$ is called divergence-free or incompressible if $\nabla\cdot\mathbf{u}=0$. We define the constraint distance $\ell_{\text{div}}$ as the $\ell^{1}$-norm of the divergence and demonstrate this constraint with 2D Kolmogorov flows [14, 10, 56]. We represent the trajectory of the 2D velocity, discretized in space-time, as a two-channel (for both velocity components) image $\mathbf{x}$ with the states appended sequentially. We used jax-cfd [44] to generate trajectories of eight $64\times 64$ states and appended them in a $2\times 4$ pattern to create $128\times 256$ images.
@@
To demonstrate a symmetry constraint, we consider images $\mathbf{x}$ that are periodic tilings of a unit cell. This type of symmetry appears in materials science, such as when constructing metamaterials out of unit cells [52]. We use a distance function $\ell_{\text{periodic}}$ that compares all pairs of tiles in the image and computes the average $\ell^{1}$-norm of their differences. We created a dataset of $64\times 64$ images (composed of $32\times 32$ unit cells tiled in a $2\times 2$ fashion) using the code of Ogren et al. [52].
@@
Generative models are known to sometimes generate the incorrect number of objects [53]. We formulate a differentiable count constraint by relying on a CNN to estimate the count of a particular object in an image $\mathbf{x}$. Note that using a neural network leads to a non-analytical and highly non-convex constraint. Letting $f_{\text{CNN}}:\mathbb{R}^{d}\to\mathbb{R}$ be the trained counting CNN, we use the distance function $\ell_{\text{count}}(\mathbf{x}):=\left\lvert f_{\text{CNN}}(\mathbf{x})-\bar{c}\right\rvert$ for a target count $\bar{c}$. The constraint dataset consists of $128\times 128$ simulated images of exactly eight (8) radio galaxies with background noise [18].
@@ 4.1 Improved constraint satisfaction and training efficiency
First and foremost, we verify that our approach leads to better constraint satisfaction than a vanilla diffusion model (DM). We evaluate constraint satisfaction by computing the average constraint distance of generated samples. We note that since the constraint distance is non-negative, an average constraint distance of 00 implies that the constraint is satisfied almost surely.
For each constraint, we trained a NAMM on the corresponding dataset and then trained an MDM on the pushforward of the dataset through the learned $\mathbf{g}_{\phi}$. We show results from a finetuned NAMM, but as we discuss in Sec. 4.3, finetuning is often not necessary. For a baseline, we trained a DM on the original dataset. Fig. 3 highlights that our MDM samples inverted through $\mathbf{f}_{\psi}$ are closer to the constraint set than DM samples. Our samples and the DM samples are visually indistinguishable, yet their constraint distances can be drastically different. For the total brightness, 1D Burgers’, divergence-free, and periodic constraints, there is a significant gap between our distribution of constraint distances and the baseline distribution. The gap is smaller, although still notable, for the count constraint, which may be due to difficulty of optimizing a highly non-convex constraint.
Furthermore, our approach achieves better constraint satisfaction in less training time. In Fig. 4, we plot constraint satisfaction as a function of compute time, comparing our approach (with and without finetuning) to the DM. Accounting for the time it takes to train the NAMM, our NAMM-based MDM achieves much lower constraint distances than the DM for the three physics-based constraints and the periodic constraint, often reaching a level that the DM struggles to achieve. For the count constraint, we find that finetuning is essential for improving constraint satisfaction, and it is more time-efficient to finetune the inverse map than to continue training the MDM.
@@ 4.2 Solving constrained inverse problems with mirror DPS
Many methods have been proposed to use a pretrained diffusion model to sample images from the posterior distribution $p(\mathbf{x}\mid\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{x})p(\mathbf{x})$ [16, 32, 63, 39, 40, 59], given measurements $\mathbf{y}\in\mathbb{R}^{m}$ and a diffusion-model prior $p(\mathbf{x})$. One of the most popular methods is diffusion posterior sampling (DPS) [17]. To adapt DPS for the mirror space, we simply evaluate the measurement likelihood on inverted mirror images $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})$ instead of on images $\mathbf{x}$ in the original space.
We demonstrate mirror DPS on data assimilation, an inverse problem that aims to recover the hidden state of a dynamical system given imperfect observations of the state. In Fig. 5, we show results for data assimilation of a 1D Burgers’ system and a divergence-free Kolmogorov flow given a few noisy state observations, which can be essentially formulated as a denoise-and-inpaint problem. For each case, we used mirror DPS with the corresponding NAMM-based MDM. We include two baselines: (1) vanilla DPS with the DM and (2) constraint-guided DPS (CG-DPS) with the DM. The latter incorporates the constraint distance as an additional likelihood term. As Fig. 5 shows, our approach leads to notably less constraint violation (i.e., less deviation from the PDE or less divergence) than both baselines. Appendix A.3 shows that our method consistently outperforms the baselines for different measurement-likelihood and constraint-guidance weights used in DPS and CG-DPS.
@@ 4.3 Ablation studies
Tab. 1 shows the improvement in constraint satisfaction after finetuning $\mathbf{f}_{\psi}$. We also verify that the resulting image distribution does not become drastically different from the true distribution by using average cosine distance between PCA components as a proxy for the similarity of the true and generated distributions. We find that finetuning does not significantly change the distribution-matching accuracy and in some cases improves it while improving constraint satisfaction.
@@
Before FT Constr. dist. ($\downarrow$) $0.49\pm 0.38$ $0.09\pm 0.01$ $2.52\pm 0.20$ $0.05\pm 0.06$ $0.73\pm 0.53$
PCA dist. ($\downarrow$) $\mathbf{0.99}$ $\mathbf{0.98}$ $0.98$ $\mathbf{1.03}$ $1.00$
After FT Constr. dist. ($\downarrow$) $\mathbf{0.49\pm 0.30}$ $\mathbf{0.04\pm 0.00}$ $\mathbf{1.42\pm 0.32}$ $\mathbf{0.04\pm 0.06}$ $\mathbf{0.36\pm 0.26}$
PCA dist. ($\downarrow$) $1.01$ $1.00$ $\mathbf{0.97}$ $\mathbf{1.03}$ $\mathbf{0.96}$
Table 1: Effect of finetuning. Constr. dist. (mean $\pm$ std. dev. of $128$ inverted MDM samples) is computed as $10\lambda_{\text{constr}}\ell$. PCA dist. is the average cosine distance between the learned distribution’s principal components and the true distribution’s principal components. PCA was done with $50$ components and approximated with $128$ inverted MDM samples and all the samples from the true test data. PCA dist. is a rough proxy that we use to confirm that finetuning does not significantly change the MDM’s distribution-matching accuracy while improving the constraint distance. For relatively simple constraints (e.g., total brightness and periodic), finetuning is not critical.
Constraint loss hyperparameters
@@
There are two hyperparameters for the constraint loss in Eq. 1: $\sigma_{\text{max}}$ determines how much noise to add to samples from the mirror distribution, and $\lambda_{\text{constr}}$ is the weight of the constraint loss. Intuitively, a higher $\sigma_{\text{max}}$ means that the inverse map $\mathbf{f}_{\psi}$ must map larger regions of $\mathbb{R}^{d}$ back to the constraint set, making its learning objective more challenging. We would expect the forward map $\mathbf{g}_{\phi}$ to cooperate by maintaining a reasonable SNR in the noisy mirror distributions. For example, Fig. 6 shows how increasing $\sigma_{\text{max}}$ causes $\mathbf{g}_{\phi}(\mathbf{x})$ for $\mathbf{x}\sim p_{\text{data}}$ to have larger magnitudes so that the added noise will not hide the signal. However, setting $\sigma_{\text{max}}$ too high can worsen constraint satisfaction, perhaps due to the challenge of finding an inverse map that effectively maps the larger sequence of noisy distributions back to the constraint set. On the flip side, setting $\sigma_{\text{max}}$ too low can also worsen constraint satisfaction because of poor robustness of $\mathbf{f}_{\psi}$. Meanwhile, increasing $\lambda_{\text{constr}}$ for the same $\sigma_{\text{max}}$ leads to lower constraint distance, although there is a tradeoff between constraint distance and cycle-consistency inherent in the NAMM objective (Eq. 1).
@@
Figure 6: Effect of $\sigma_{\text{max}}$ and $\lambda_{\text{constr}}$, demonstrated with 1D Burgers’. First row: as $\sigma_{\text{max}}$ increases (keeping $\lambda_{\text{constr}}=1.0$), the mirror image $\mathbf{g}_{\phi}(\mathbf{x})$ for $\mathbf{x}\sim p_{\text{data}}$ increases in magnitude to maintain a similar SNR. Histograms show constraint distances of $128$ inverted MDM samples, normalized to have a maximum of 1 across the samples from all three settings. Decreasing $\sigma_{\text{max}}$ from $0.5$ to $0.1$ improves the constraint distances, but further lowering it to $0.001$ causes the constraint distances to go back up. This indicates a tradeoff between robustness and performance of $\mathbf{f}_{\psi}$. Second row: as $\lambda_{\text{constr}}$ increases (keeping $\sigma_{\text{max}}=0.1$), $\mathbf{g}_{\phi}(\mathbf{x})$ does not change as much as when increasing $\sigma_{\text{max}}$, but the constraint distances decrease (with a tradeoff in cycle consistency). For all three settings, the same number of NAMM and MDM epochs was used as in Fig. 3 but without finetuning.
@@
Fig. 7 compares parameterizing $\mathbf{g}_{\phi}$ as the gradient of an ICNN versus as a ResNet-based CNN. We find that the ICNN approach has practical benefits in addition to being theoretically justified. For the 1D Burgers’ and divergence-free constraints, we demonstrate how the mirror space changes when instead parameterizing the forward map as a ResNet-based CNN. The mirror space becomes less regularized, leading to worse constraint satisfaction of the MDM. A possible explanation is that the MDM struggles to learn a less-regularized mirror space.
Figure 7: Parameterizing $\mathbf{g}_{\phi}$ as gradient of ICNN vs. as ResNet-based CNN. Both types of mirror maps preserve some visual similarity, but using an ICNN better regularizes the mirror space. The ResNet mirror images include irregularities, such as the patch circled in yellow in the divergence-free example. The histograms show the normalized constraint distances $\bar{\ell}$ of $128$ inverted MDM samples (without finetuning). The rightmost histograms include the DM results from Fig. 3 for comparison. As the histograms show, an ICNN leads to better constraint satisfaction with fewer outliers. We followed the number of NAMM and MDM epochs used in Fig. 3 for both constraints without finetuning. We found that even with finetuning, a ResNet-based forward map still leads to inverted MDM samples of worse constraint satisfaction or noticeably worse visual quality.

### v2

@@ 3.2 Learning the mirror diffusion model
Similarly to Liu et al. (2024), we train an MDM on the mirror distribution ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$ and map its samples to the constrained space through $\mathbf{f}_{\psi}$. In particular, we train a score model $\mathbf{s}_{\theta}$ with the following denoising score matching objective in the learned mirror space (defined as the range of $\mathbf{g}_{\phi}$):
@@ 3.3 Finetuning the inverse mirror map
The inverse map $\mathbf{f}_{\psi}$ is trained with samples from the noisy mirror distributions in Equation 2, but we ultimately wish to evaluate $\mathbf{f}_{\psi}$ with samples from the MDM. To reduce the distribution shift, it may be helpful to finetune $\mathbf{f}_{\psi}$ with MDM samples. We generate a training dataset of samples $\tilde{\mathbf{x}}$ from the MDM and then finetune the inverse map to deal with such samples specifically. In the following finetuning objective, we replace $\tilde{\mathbf{x}}\sim{(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$ with $\tilde{\mathbf{x}}\sim p_{\theta}$, where $p_{\theta}$ is the distribution of MDM samples in the mirror space:
@@
Finetuning essentially tailors $\mathbf{f}_{\psi}$ to the MDM. The original objective assumes that the MDM will sample Gaussian-perturbed images from the mirror distribution, but in reality it samples from a slightly different distribution. As the ablation study in Section 4.3 shows, finetuning is not an essential component of the method; we suggest it as an optional step for when it is critical to optimize the constraint distance metric.
@@ 4 Results
We present experiments with constraints ranging from physics-based to semantic. For the considered examples, our method achieves from $38\%$ to as much as $96\%$ improvement in constraint satisfaction upon a vanilla DM trained on the same data (see Table 1). Appendices D.1 and F provide implementation and constraint details, respectively. The following paragraphs introduce the demonstrated constraints $\ell$. For each we consider an image dataset for which the constraint is physically meaningful.
@@
In astronomical imaging, even if a source’s structure is unknown a priori, its total brightness, or total flux, is often well constrained (EHTC, 2019). We define $\ell_{\text{flux}}(\mathbf{x})$ as the absolute difference between $\sum_{i=1}^{d}\mathbf{x}_{i}$ and the true total brightness. We demonstrate with a dataset of $64\times 64$ images of black-hole simulations (Wong et al., 2022) whose pixel values sum to $120$. While this constraint is a simple warmup example, generic diffusion models perform surprisingly poorly on it.
@@
We consider Burgers’ equation (Bateman, 1915; Burgers, 1948) for a 1D viscous fluid, representing the discretized solution as an $n_{x}\times n_{t}$ image $\mathbf{x}$, where $n_{x}$ and $n_{t}$ are the numbers of grid points in space and time, respectively. The distance $\ell_{\text{burgers}}(\mathbf{x})$ compares each 1D state in the image to the PDE solver’s output given the previous state (based on Crank-Nicolson time-discretization (Crank & Nicolson, 1947; Kidger, 2022)). The dataset consists of $64\times 64$ images of Crank-Nicolson solutions with Gaussian random fields as initial conditions.
@@
A time-dependent 2D velocity field $\mathbf{u}=\mathbf{u}(x,y,t)$ is called divergence-free or incompressible if $\nabla\cdot\mathbf{u}=0$. We define the constraint distance $\ell_{\text{div}}$ as the $\ell^{1}$-norm of the divergence and demonstrate this constraint with 2D Kolmogorov flows (Chandler & Kerswell, 2013; Boffetta & Ecke, 2012; Rozet & Louppe, 2024). We represent the trajectory of the 2D velocity, discretized in space-time, as a two-channel (for both velocity components) image $\mathbf{x}$ with the states appended sequentially. We used jax-cfd (Kochkov et al., 2021) to generate trajectories of eight $64\times 64$ states and appended them in a $2\times 4$ pattern to create $128\times 256$ images.
@@
We consider images $\mathbf{x}$ that are periodic tilings of a unit cell. This type of symmetry appears in materials science, such as when constructing metamaterials out of unit cells (Ogren et al., 2024). We use a distance function $\ell_{\text{periodic}}$ that compares all pairs of tiles in the image and computes the average $\ell^{1}$-norm of their differences. We created a dataset of $64\times 64$ images (composed of $32\times 32$ unit cells tiled in a $2\times 2$ fashion) using code from Ogren et al. (2024).
@@
Generative models can sometimes generate an incorrect number of objects (Paiss et al., 2023). We formulate a differentiable count constraint by relying on a CNN to estimate the count of a particular object in an image $\mathbf{x}$. Note that using a neural network leads to a non-analytical and highly non-convex constraint. Letting $f_{\text{CNN}}:\mathbb{R}^{d}\to\mathbb{R}$ be the trained counting CNN, we use the distance function $\ell_{\text{count}}(\mathbf{x}):=\left\lvert f_{\text{CNN}}(\mathbf{x})-\bar{c}\right\rvert$ for a target count $\bar{c}$. The dataset consists of $128\times 128$ simulated images of exactly eight (8) radio galaxies with background noise (Connor et al., 2022).
@@ 4.1 Improved constraint satisfaction and training efficiency
First and foremost, we verify that our approach leads to better constraint satisfaction than a vanilla diffusion model (DM). We evaluate constraint satisfaction by computing the average constraint distance of generated samples. Since the constraint distance is non-negative, an average constraint distance of 00 implies that the constraint is satisfied almost surely.
For each constraint, we trained a NAMM on the corresponding dataset and then trained an MDM on the pushforward of the dataset through the learned $\mathbf{g}_{\phi}$. We show results from a finetuned NAMM, but as shown in Section 4.3, finetuning is often not necessary. The baseline DM was trained on the original dataset. Figure 3 highlights that MDM samples inverted through $\mathbf{f}_{\psi}$ are much closer to the constraint set than DM samples despite being visually indistinguishable. For the total brightness, 1D Burgers’, divergence-free, and periodic constraints, there is a significant gap between our distribution of constraint distances and the baseline’s. The gap is smaller for the count constraint, which may be due to difficulties in identifying and learning the mirror map for a highly non-convex constraint.
Furthermore, our approach achieves better constraint satisfaction in less training time. In Figure 4, we plot constraint satisfaction as a function of compute time, comparing our approach (with and without finetuning) to the DM. Accounting for the time it takes to train the NAMM, our MDM achieves much lower constraint distances than the DM for the three physics-based constraints and the periodic constraint, often reaching a level that the DM struggles to achieve. For the count constraint, we find that finetuning is essential for improving constraint satisfaction, and it is more time-efficient to finetune the inverse map than to continue training the MDM.
@@ 4.2 Solving constrained inverse problems with mirror DPS
Many methods have been proposed to use a pretrained diffusion model to sample images from the posterior distribution $p(\mathbf{x}\mid\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{x})p(\mathbf{x})$ (Choi et al., 2021; Graikos et al., 2022; Song et al., 2022; Jalal et al., 2021; Kawar et al., 2022; Song et al., 2023), given measurements $\mathbf{y}\in\mathbb{R}^{m}$ and a diffusion-model prior $p(\mathbf{x})$. One of the most popular methods is diffusion posterior sampling (DPS) (Chung et al., 2022). To adapt DPS for the mirror space, we simply evaluate the measurement likelihood on inverted mirror images $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})$ instead of on images $\mathbf{x}$ in the original space.
We demonstrate mirror DPS on data assimilation, an inverse problem that aims to recover the hidden state of a dynamical system given imperfect observations of the state. In Figure 5, we show results for data assimilation of a 1D Burgers’ system and a divergence-free Kolmogorov flow given a few noisy state observations, which can be essentially formulated as a denoise-and-inpaint problem. For each case, we used mirror DPS with the corresponding NAMM-based MDM. We include two baselines: (1) vanilla DPS with the DM and (2) constraint-guided DPS (CG-DPS) with the DM. The latter incorporates the constraint distance as an additional likelihood term. As Figure 5 shows, our approach leads to notably less constraint violation (i.e., less deviation from the PDE or less divergence) than both baselines. Appendix D.3 shows that our method consistently outperforms the baselines for different measurement-likelihood and constraint-guidance weights used in DPS and CG-DPS.
@@ 4.3 Ablation studies
Table 1 shows the improvement in constraint satisfaction after finetuning $\mathbf{f}_{\psi}$ while also verifying that the generated distribution stays close to the true data distribution. We use maximum mean discrepancy (MMD) (Gretton et al., 2012) and Kernel Inception Distance (KID) (Bińkowski et al., 2018) as measures of distance between distributions. MMD evaluates distance in a feature space defined by a Gaussian kernel, and KID uses Inception v3 (Szegedy et al., 2015) features (see Appendix E for details). Finetuning does not notably change the distribution-matching accuracy of the MDM and in some cases improves it while improving constraint satisfaction. Compared to a vanilla DM, our approach before and after finetuning does not lead to significantly different MMD and even gives better KID while significantly improving constraint distance.
@@
Ours w/o FT CD ($\downarrow$) $0.57\pm 0.48$ $0.09\pm 0.01$ $2.51\pm 0.21$ $0.04\pm 0.05$ $0.71\pm 0.53$
MMD ($\downarrow$) $0.0957\pm 0.0130$ $0.1225\pm 0.0096$ $0.0664\pm 0.0027$ $0.0760\pm 0.0058$ $0.1716\pm 0.0062$
KID ($\downarrow$) $\mathbf{0.0026}\pm 0.0007$ $0.0040\pm 0.0004$ $\mathbf{0.0035}\pm 0.0004$ $0.0018\pm 0.0005$ $\mathbf{0.0367}\pm 0.0010$
Ours CD ($\downarrow$) $\mathbf{0.49}\pm 0.35$ $\mathbf{0.04}\pm 0.01$ $\mathbf{1.57}\pm 0.31$ $\mathbf{0.04}\pm 0.06$ $\mathbf{0.35}\pm 0.27$
MMD ($\downarrow$) $0.1023\pm 0.0131$ $0.1291\pm 0.0096$ $0.0781\pm 0.0023$ $0.0758\pm 0.0058$ $0.1978\pm 0.0062$
KID ($\downarrow$) $0.0027\pm 0.0008$ $\mathbf{0.0022}\pm 0.0004$ $0.0058\pm 0.0006$ $\mathbf{0.0014}\pm 0.0005$ $0.0570\pm 0.0014$
Baseline CD ($\downarrow$) $2.20\pm 1.61$ $0.45\pm 0.03$ $6.96\pm 0.27$ $1.04\pm 0.08$ $0.56\pm 0.44$
MMD ($\downarrow$) $\mathbf{0.0956}\pm 0.0099$ $\mathbf{0.0621}\pm 0.0091$ $\mathbf{0.0595}\pm 0.0026$ $\mathbf{0.0533}\pm 0.0043$ $\mathbf{0.1276}\pm 0.0066$
KID ($\downarrow$) $0.0462\pm 0.0016$ $0.2308\pm 0.0028$ $0.0053\pm 0.0007$ $0.0064\pm 0.0004$ $0.1084\pm 0.0015$
Table 1: Effect of finetuning. Constr. dist. (CD) $=100\lambda_{\text{constr}}\ell$. The improvements in mean CD are (left to right, comparing “Ours” to “Baseline”): $78\%$, $91\%$, $77\%$, $96\%$, $38\%$ for five problems. For all metrics, mean $\pm$ std. dev. is estimated with $10000$ samples. In terms of MMD/KID, finetuning does not significantly impact distribution-matching accuracy but improves constraint distance. DM baseline results are shown for comparison. According to MMD, the baseline gives better distribution-matching accuracy; according to KID, our approach captures the true data distribution better.
Constraint loss
@@
There are two hyperparameters for the constraint loss in Equation 1: $\sigma_{\text{max}}$ determines how much noise to add to samples from the mirror distribution, and $\lambda_{\text{constr}}$ is the weight of the constraint loss. Intuitively, a higher $\sigma_{\text{max}}$ means that the inverse map $\mathbf{f}_{\psi}$ must map larger regions of $\mathbb{R}^{d}$ back to the constraint set, making its learning objective more challenging. We would expect $\mathbf{g}_{\phi}$ to cooperate by maintaining a reasonable SNR in the noisy mirror distributions. Figure 6 shows how increasing $\sigma_{\text{max}}$ causes $\mathbf{g}_{\phi}(\mathbf{x})$ for $\mathbf{x}\sim p_{\text{data}}$ to have larger magnitudes so that the added noise will not hide the signal. However, setting $\sigma_{\text{max}}$ too high can worsen constraint satisfaction, perhaps due to the challenge of mapping a larger region of $\mathbb{R}^{d}$ back to the constraint set. On the flip side, setting $\sigma_{\text{max}}$ too low can worsen constraint satisfaction because of poor robustness of $\mathbf{f}_{\psi}$. Meanwhile, increasing $\lambda_{\text{constr}}$ for the same $\sigma_{\text{max}}$ leads to lower constraint distance, although there is a tradeoff between constraint distance and cycle-consistency inherent in the NAMM objective (Equation 1).
@@
Figure 6: Effect of $\sigma_{\text{max}}$ and $\lambda_{\text{constr}}$, demonstrated with 1D Burgers’. First row: as $\sigma_{\text{max}}$ increases (keeping $\lambda_{\text{constr}}=1.0$), the mirror image $\mathbf{g}_{\phi}(\mathbf{x})$ for $\mathbf{x}\sim p_{\text{data}}$ increases in magnitude to maintain a similar SNR. Histograms show constraint distances of $128$ inverted MDM samples, normalized to have a maximum of $1$ across samples from all three settings. Decreasing $\sigma_{\text{max}}$ from $0.5$ to $0.1$ improves the constraint distances, but further lowering $\sigma_{\text{max}}$ to $0.001$ causes them to go back up. This indicates a tradeoff between robustness and performance of $\mathbf{f}_{\psi}$. Second row: as $\lambda_{\text{constr}}$ increases (keeping $\sigma_{\text{max}}=0.1$), $\mathbf{g}_{\phi}(\mathbf{x})$ does not change as much as when increasing $\sigma_{\text{max}}$, but the constraint distances decrease (with a tradeoff in cycle consistency). For all three settings, the same number of NAMM and MDM epochs was used as in Fig. 3 but without finetuning.
@@
Figure 7 compares parameterizing $\mathbf{g}_{\phi}$ as the gradient of an ICNN versus as a ResNet-based CNN. We demonstrate how the mirror space changes when parameterizing the forward map as a ResNet-based CNN. The mirror space becomes less regularized, leading to worse constraint satisfaction of the MDM, perhaps because the MDM struggles to learn a less-regularized mirror space. Thus, even when the constraint is non-convex and there are no theoretical reasons to use an ICNN, it may still be practically favorable.
Figure 7: Architecture of $\mathbf{g}_{\phi}$: gradient of ICNN vs. ResNet-based CNN. Both approaches preserve visual structure in the mirror space, but the ResNet causes irregularities, such as the patch circled in yellow. The histograms show the normalized constraint distances $\bar{\ell}$ of $128$ inverted MDM samples (without finetuning). DM histograms from Fig. 3 are shown for comparison. An ICNN leads to better constraint satisfaction with fewer outliers. We trained the NAMM and MDM for the same number of epochs as in Fig. 3 without finetuning. We found that even with finetuning, a ResNet-based forward map leads to worse constraint satisfaction or noticeably worse visual quality of generated samples.